# Deliberative XAI: How Explanations Impact Understanding and Decision-Making of AI Novices in Collective and Individual Settings

### v1

@@ 4. Results
G3: I think that’s a bit weird. […] Because if she can speak English very well and has the specialist knowledge that our IT sector needs… (analyzing: discussing)
G4: She gets two minuses for living in [a specific city district777Information was omitted to adhere to anonymization policy and will be re-inserted for the final version.]. People who live in [this district], or in a district that is not well regarded, have totally negative… (analyzing: explaining)
G1: Yes, and here you have it in writing, I’ll have to look at that too. (sharing: swapping)
Framed in the context of mechanisms that drive collaborative success, G4 first draws the group’s attention to an explanation they deem relevant, describing the algorithm’s treatment of two different cases of job-seekers. G4 reads the information to the group, which prompts G2 and G3 to contribute their own perspectives, ”increasing engagement” while also ”cueing their prior knowledge”. As a result, G1 is also engaged by the information and articulates interest in looking at it. In combination, the interaction has established ”common ground” by sharing information the group can use for further collaboration. Interactions like these were common in Groups G and H (job counselors and trainers), indicating that the explanations’ format and content suited these compositions.
Case 2: Discussion of solutions leads to correction. Group B consisted of both staff members and volunteers of a civil society organization. The group interaction was mostly driven by one participant who requested and shared most of the information and thus acted as a guide and informer, potentially driven by her role as deputy manager. While this led to a more one-sided dynamic, the other participants later appreciated outsourcing part of the understanding process to B3. In the excerpt below, B3 proposes a solution for task 2 (chance of employment that the system assigns Harald G.; group low, medium, or high) that is then discussed:
B3: I chose “group low” for task 2. (sharing: proposing solutions)
B4: I put him [Harald G.] in ”group medium”. (sharing: proposing solutions)
B1: Why not medium, B3? (analyzing: discussing) […]
B3: With 49 years he has one minus. And the knee problems could be a “health impairment”. (analyzing: explaining)
B2: But if he has a job where he sits, that doesn’t matter. (analyzing: discussing)
B3: But I don’t see that this is considered anywhere. And I think the fact that he hasn’t been employed much in the last 4 years has a very negative effect. What information do you have that speaks against that? (combining: correcting; analyzing: discussing)
B1: Because with 50, he would still be in ”group medium”. […] Then he’s male and from [this country888Information was omitted to adhere to anonymization policy and will be re-inserted for the final version.]. (analyzing: explaining)
B3: Male is only 0, the citizenship is also only 0. That doesn’t make up for anything. (combining: correcting)
[B1 agrees] […]
B4: Then I’ll correct it to low again. (combining: merging)
The group members in this situation ”negotiate multiple perspectives” on the correct solution, notice discrepancies in their ”common ground,” and ”re-expose” themselves to information to ”correct errors” in their knowledge. B3 led the interaction and corrected other participants repeatedly, which could have caused adverse reactions from them (e.g., due to being ”evaluated”). However, B2 and B4 in fact later articulated appreciation for this interaction, stating that ”B3 filtered that out well” (B2) and that ”it was helpful that you [B3] always read it out and corrected it straight away” (B4). Correction was therefore seen as constructive to the discussion, possibly because of B3’s leading role in the organization and because participants already knew each other.
Case 3: Attempted error-correction fails to create common ground. Group E was composed of people who were job-seeking at the time of the interview or in the past. Members of this group were not familiar with each other. Interaction was mostly fragmented and participants rarely shared or discussed information. In the excerpt below, participants were asked about their responses to task 4 (whether Harald G. is able to contest the system’s decision):
E2: I say you can always object. Even if you’re locked up, you can go to court somewhere. (sharing: proposing solutions)
E3: Yes, I think so too. (sharing: proposing solutions)
E1: Well, I’ve already read that [the explanation], and it says that you can’t appeal against it. But it says that you could if a legal basis were first created, which doesn’t exist at the moment. (combining: correcting; analyzing: explaining)
Investigator: Okay. Is there anything you [E2, E3] would like to change now, based on this information?
[E2 & E3 indicate that they do not]
Investigator: You would say that he can contest the decision?
[E2 & E3 agree]
Despite E1’s reference to an explanation and their attempt to ”correct errors” in what E2 and E3 believe to be the solution, the interaction fails to establish ”common ground” or ”negotiate multiple perspectives”, as the different viewpoints are neither elaborated nor discussed collectively. Possibly, E2 and E3 did not believe that E1 or the explanation were correct, or they felt embarrassed for being corrected. Further, E1 had no particular motivation to persuade the other participants of the correct solution, as E1 later self-described as ”introverted” and stated that they usually preferred individual over group settings. Comparable mechanisms unfolded only in Group F, where participants tended to read explanations themselves or disengage rather than interact with other participants. This might have had multiple reasons: First, both Groups E and F indicated a low interest in the topic of labor; second, participants in both groups had trouble understanding the explanations due to language barriers and the heavy reliance on text; and third, participants in both groups were unknown to each other. In combination, this might have created a group dynamic that was not constructive, i.e., neither argumentative (Mercier and Sperber, 2011), solution-oriented (Niculae and Danescu-Niculescu-Mizil, 2016), or familiar (Johnson and Johnson, 1985), which disrupted collaborative interaction with the explanations as well as discussion processes.
Summary. The conversation excerpts in cases 1 and 2 demonstrate that the explanation design supported various collaborative interaction processes. These processes led participants to create common ground and find a task solution through error correction, which are markers of ”collaborative success” (Nokes-Malach et al., 2015). In contrast, case 3 demonstrates that the explanations did not always achieve their purpose of facilitating interaction, highlighting that the interplay of explanation format and group dynamics must be considered in the design of explanations in group settings.
4.2.2. Explanations in group settings allow for informed discussion of pivotal topics and opinion formation about system deployment.
For most participants, deciding about the employment prediction algorithm’s deployment was a clear choice: 7 out of 8 groups voted ”No”, and a majority of participants also voted ”No” in their individual decisions (Table 4). Confidence in these decisions increased strongly after the explanation phase (Table 3), with most participants indicating that they now felt better informed. Comparing single interviews with focus groups, two effects were notable: First, a third of focus group participants changed their vote in the second self-report, and second, participants in single interviews stated repeatedly that they would have found a discussion helpful.
This is further underscored by participants’ feedback on the focus groups: Most felt included in the discussions (26 of 31) and were influenced in their decisions by them (18 of 31). Specifically, participants appreciated the group size, which allowed for balanced contributions and productive discussion (C1, D1, E2), and valued the exchange of perspectives: ”Everyone understands a little differently and sees a different focus, I think it’s really important to talk about that.” (H5). We see this as evidence that the explanations were suitable for the collaborative setting and supported group deliberation. This section thus aims to describe the synergy effects that emerge when combining explanations with group discussion and how this affected participants’ perceptions. In the following, we describe three conversation excerpts in which explanations lead groups to discuss their core concerns and beliefs.

### v2

@@ 4. Results
G3 I think that’s a bit weird. […] Because if she can speak English very well and has the specialist knowledge that our IT sector needs…
G4 She even gets two minuses for living in Favoriten [a city district]. […]
G1 Yes, and here you have it in writing, I’ll have to look at that too.
A later excerpt in Section 4.2.3 further shows that explanations also led participants to ‘negotiate multiple perspectives’. Addressing both sides of collaboration, cognitive and social, here is an important goal when supporting group understanding. It can be assumed that when explanations succeed in doing both, they can provide participants with more comprehensive and more complex information than individual settings. In line with this, previous work in HCI has found that group interaction boosts task performance compared to individual settings (Karadzhov et al., 2023). However, in our study, individuals surprisingly performed better in the study tasks than groups, as discussed in Section 4.1.3. Even so, we argue that the set of interactions (shared understanding) enabled through the combination of our explanation design and the group setting presents important pathways to help AI novices understand algorithmic systems. These interactions can be especially useful when group members have different domain expertise and information needs, as they can use complementary knowledge, memory, and perspectives to make sense of information. At the same time, as noted earlier, these interactions are partly dependent on the group dynamic. Positive interactions like those described were especially frequent in Groups G and H (job counselors and trainers), where participants knew and trusted each other. In contrast, the next subsection describes instances where groups encountered challenges in understanding, illustrating the importance of social mechanisms.
4.1.2. Groups’ drawbacks: Process loss and susceptibility to social dynamics
In some of the focus groups, participants lost track of information, forgot their train of thought, or abandoned understanding altogether. We summarize these effects under the term impeded understanding and its final result as abandoned understanding. We found that impeded understanding occurred due to explanation design flaws and co-occurred with adverse social dynamics, resulting in ‘process loss’ (groups falling short of their potential performance (Kerr and Tindale, 2004)). For some participants, the benefits of the explanations’ modular structure turned into disadvantages when it hampered them in navigating and retrieving information. Such impeded interactions included cumbersome information uptake, being overwhelmed by information, and relying on intuition over information. Further, for some participants, the group setting contributed to these impediments. For example, participant B3 stated that “For me, it doesn’t make sense to […] split up [the explanations], and everyone reads a part, that’s actually not enough.” As before, these impediments can be connected to cognitive mechanisms of ‘collaborative failure’ (Nokes-Malach et al., 2015). When groups had difficulties in interacting with the explanations, they also incurred ‘memory coordination cost’ (increased cognitive load) and ‘retrieval strategy disruption’ (losing train of thought). We illustrate these mechanisms with an exchange from Group G. Although this group was composed of participants with university education, it did not succeed in calculating the employment chance for study task 2, in contrast to single interview participants with the same education.
G4 We can go through the features briefly. Where is the piece of paper with this terrible matrix? […]
G2 I still don’t understand which value to put. To calculate it, I need an exact value for the weighting.
G1 You can calculate it with this. The apprenticeship has 52%. I believe that he [Harald] has over 25%.
G3 Yes, definitely, I mean, roughly speaking…
G1 She [Shifteh] has over 30%. And she also has 2 minuses […] and a plus.
G3 That’s also how I estimated it. […]
G2 But how do you calculate it? […] And why are there differences between the general weighting and the exact calculation? That doesn’t click for me right now.
Here, the levels of detail in the explanations acted against participants’ understanding by obscuring the actual feature weights, which were only accessible in level 2 of system details. While improved navigation might solve this issue, it also shows the difficulty of simplifying information about AI systems without omitting key aspects. In the intention to provide an easier reading of the feature weights, which was an advantage in other cases, the explanations’ clarity was reduced, and information was obscured. Addressing all information needs of AI novices (Schmude et al., 2024) thus leads to problems with information overload, as observed in previous work on ‘white-box’ explanations (Cheng et al., 2019). However, impeded understanding alone did not mean that collaboration failed; rather, it depended on how groups dealt with these issues. Here, the key aspects were group cohesion (Kerr and Tindale, 2004) and constructiveness (Niculae and Danescu-Niculescu-Mizil, 2016). When the group dynamic was unfamiliar, it gave room to negative social mechanisms, such as ‘social loafing’ (group loses motivation) and ‘fear of evaluation’ (being criticized by others), and participants began to abandon understanding. This suggests that when interaction between participants stopped, interaction with the explanations stopped as well. Limiting these negative dynamics and promoting positive ones must be a goal of both explanation design and setting.
Notably, these adverse social dynamics occurred most often in Groups E and F, which were composed of job-seekers. Participants had trouble engaging with the explanations and abandoned interactions and understanding by saying: “Probably [you can solve it] with that, but I don’t know, I’m too stupid for that.” (E3) or “I don’t know what I should say. Everything has already been said.” (F2). Here, two things failed: The explanations failed to make crucial information accessible, and the group setting failed to uplift members who were discouraged. Interactions that offset this discouragement, such as locating information together and outsourcing, were not realized in Groups E and F. We thus propose to use co-design approaches to make explanations viable for decision subjects, as has been done with public servants (Weitz et al., 2024). Further, XAI should employ methods that create a productive social dynamic, which we identify as the second key aspect to support ‘collaborative success’ and shared understanding (Figure 5).
4.1.3. Why individuals performed better in the study tasks but still felt the absence of collaboration.
Despite the different advantages and disadvantages that group and individual settings offer in learning contexts (Section 2.3.1), these differences are rarely examined empirically in XAI888In one study that made this comparison, participants met in a chatroom and decided in collaboration with the AI and not about it (Chiang et al., 2023). We address this gap by comparing in-person focus groups and single interviews to examine whether the social setting impacted participants’ understanding. We use a triangulation999Using multiple methods of data collection (here: self-reports, task performance, articulations/actions) to explore a phenomenon (Carter et al., 2014). approach by investigating participants’ understanding with respect to three aspects: interactions with the explanations, task performance, and self-reported understanding.
Overall, the explanation afforded its core functionality to both individuals and groups. Interestingly, participants in single interviews tended to request the same number or more explanations than focus groups (Figure 6), despite having seemingly less working memory. S5 described: “Well, maybe [it was overwhelming] at the very beginning […] But I then realized that I could get through it to some extent.” Single interview participants also regularly consumed explanations in bulk, i.e., read through the whole of an explanation category rapidly. This interaction was nearly nonexistent in focus groups. However, even though participants in single interviews performed better in the study tasks, they often stated that they missed the “exchange with people, with other perspectives” (S3). S8 explained that this exchange would allow for a different form of understanding:
I think that, on your own, you can think about it very intensely and […] make up your own mind. But that’s also the disadvantage, making up your own mind. Others may have completely different thoughts and a different professional background. And that would probably have been an exciting exchange. (S8)
Participants in individual settings further performed better in the study tasks than groups (Table 3). A possible explanation is that participants in single interviews engaged differently with the study tasks, as they often calculated the exact employment chance in study task 2. None of the focus groups completed this step, but rather made educated guesses. This might be explained by the degree of focused attention the settings afforded participants. As single interviews incurred no distractions, participants could immerse themselves in the explanations.
Notably, this difference is not represented in self-reported understanding. Most participants in both settings reported unchanged understanding after the explanation phase (Table 3). Paradoxically, they verbally stated that it improved. E3 commented: “I don’t think I understood it the way you can understand it yet, but it’s definitely better than before.” And S3 explained that: “I would still say my understanding is ‘good’, but this ‘good understanding’ is much more informed now than the first superficial one.” This indicates that participants tended to judge their understanding relative to the information available, not necessarily in relation to their previous report. We describe this process as calibrating understanding. Previous research in cognitive science has documented similar effects (Keil, 2006), which were also reproduced in an XAI study on white-box explanations (Cheng et al., 2019). However, participants’ verbal reports, their feedback on the explanations (Section 4.1.4), and the calibration process itself indicate that the explanations improved understanding. Including additional measures, such as information gain, could further capture the calibration process, which is discussed in Section 5.2.
To compare individual and group interactions with the explanations, we lastly draw from the ‘six facets of understanding’ (Wiggins and McTighe, 2005). The framework describes that understanding is represented by the ability to ‘explain, interpret, apply, take perspective, empathize, and self-reflect’ with respect to a topic. The more facets are covered, the better the understanding. Seeing that individuals had a clear advantage in solving the study tasks suggests that the individual setting supported the ‘apply’ facet. In contrast, the group settings often led participants to ‘explain’ information to others and ‘interpret’ it (Section 4.1.1), and to ‘take perspective’ and ‘empathize’ through the exchange of views and experiences (Section 4.2). As explanations aim to improve understanding of a given AI system, combining both settings to cover more facets of understanding could thus be a fruitful approach. Further, explanations for individuals in particular can benefit from information covering facets usually dependent on social interaction. Our design aimed to implement this through explanations such as What chance would the system calculate for me? (interpret – making it personally relevant), and How could the system make the work of counselors easier or more difficult? (take perspective: provide multiple angles and arguments).
Based on these findings, we argue that both group and individual settings can contribute to participant understanding and should ideally be combined. In particular, focused attention can facilitate the application of information, while shared understanding and the exchange of opinions and arguments (Section 4.2) aid encouragement, reflection, and collective action. Considering this trade-off between settings can inform how explanations can be combined with social settings to cover as many understanding facets as possible.
Focus groups | |||||||||
ID | Und. I | Und. II | Change | Task Performance | ID | Und. I | Und. II | Change | Task Performance |
A1 | 2 | 2 | 0 | F1 | 2 | 4 | +2 | ||
A2 | 1 | 2 | +1 | F2 | 1 | 4 | +3 | ||
A3 | 4 | 4 | 0 | F3 | 5 | 2 | -3 | ||
A4 | 4 | 4 | 0 | F4 | 3 | 3 | 0 | ||
B1 | 4 | 2 | -2 | F5 | 4 | 4 | 0 | ||
B2 | 4 | 2 | -2 | G1 | 5 | 5 | 0 | ||
B3 | 5 | 5 | 0 | G2 | 3 | 2 | -1 | ||
B4 | 2 | 3 | +1 | G3 | 5 | 2 | -3 | ||
C1 | 4 | 4 | 0 | G4 | 5 | 4 | -1 | ||
C2 | 2 | 2 | 0 | H1 | 4 | 5 | +1 | ||
C3 | 1 | 3 | +2 | H2 | 4 | 5 | +1 | ||
D1 | 4 | 3 | -1 | H3 | 4 | 4 | 0 | ||
D2 | 4 | 3 | -1 | H4 | 4 | 5 | +1 | ||
D3 | 4 | 4 | 0 | H5 | 4 | 4 | 0 | ||
E1 | 3 | 4 | +1 | ||||||
E2 | 3 | 3 | 0 | ||||||
E3 | 4 | 4 | 0 | ||||||
Single interviews | |||||||||
ID | Und. I | Und. II | Change | Task Performance | ID | Und. I | Und. II | Change | Task Performance |
S1 | 5 | 5 | 0 | S7 | 4 | 4 | 0 | ||
S2 | 4 | 2 | -2 | S8 | 3 | 4 | +1 | ||
S3 | 4 | 4 | 0 | S9 | 5 | 5 | 0 | ||
S4 | 4 | 4 | 0 | S10 | 4 | 3 | -1 | ||
S5 | 5 | 4 | -1 | S11 | 4 | 4 | 0 | ||
S6 | 1 | 4 | +3 | S12 | 2 | 4 | +2 |
4.1.4. Reflections on the explanation design: Modularity, levels of detail, and most important information.
To examine how the explanation design was received, we asked participants for feedback on the explanations’ structure, content, style of expression, and information coverage. We report and summarize the participants’ criticisms as a basis to formulate design improvement suggestions in Section 5.
Strengths and weaknesses of the design. Positive comments described the explanations’ structure as “nicely presented” (A2, C2) and “good to get an overview” (C3, H4) while being “active and controllable” (S8). Critical comments described the information coverage as “too much” (D1, S4), and the structure as “confusing” (B1, D1) and “demanding” (D2). Participants saw the design’s strengths in its four-category structure, question-driven presentation, active selection, and information scope. However, the scope and depth of information also led to information overload and loss of overview. Further, the explanations’ many and complicated texts were described as “very difficult” (E2, F2). E2 compared the language to “letters […] from the court. I understand every single word, but I don’t understand the context.” Previous work has found that textual explanations can effectively convey information but tend to raise aversion with users (Szymanski et al., 2021; Schmude et al., 2023). However, Weitz et al. (2024) paradoxically found that users preferred textual explanations. This points to the need for further research on textual formats in XAI, like the automated adaptation of text to different difficulties.
Most helpful and influential information. Participants in focus groups stated that all explanation categories helped their understanding and influenced their decision evenly ( Figure 7), often mentioning that “all of them [are relevant]… I don’t think you can leave anything out, really” (D3). In contrast, participants in single interviews found data much less helpful and less influential, stating, e.g., that they prioritized another category in the time available. Notably, participants emphasized that two categories were central: system details and context. System details were perceived as “tangible” (S6) and “concrete” (S8), and explanations about the features and weighting were perceived as especially important: “That is the central point, the basis of the whole system.” (G4) In turn, explanations from the category context were requested the most (Figure 6). Here, participants appreciated explanations that described decision subjects’ inability to contest decisions and the system’s political background. Drawing from the concept of ‘intelligibility types’ (Lim and Dey, 2009), we argue that system details provided descriptive information to the question “What did the system do?”, while context provided normative information to the question “Why did the system do [this]?”. Future research should investigate how both information types can be integrated into explanations for AI novices.
4.1.5. Summary RQ1: Group and individual settings address different understanding facets in explainable AI
In RQ1-Explanations, we asked how a question-driven, modular explanation design supports AI novices’ understanding in groups and individual settings. We found that the explanations support both settings but differ in how understanding develops. In groups, we found that explanations facilitated interactions that produced shared understanding and involved cognitive and social mechanisms of ‘collaborative success’ (Nokes-Malach et al., 2015). In groups with a trusting social dynamic, participants tackled explanations together, which acted against discouragement. When groups had a negative social dynamic, the explanation design could become overwhelming and understanding issues were left unchecked, leading to ‘process loss’ (Kerr and Tindale, 2004) and abandoned understanding. Participants in single interviews interacted with the explanations in a more focused and self-directed manner. This had advantages for task performance and explanation engagement, which aligns with research on tutoring methods (Bloom, 1984). However, the positive effects of aggregated knowledge (Navajas et al., 2018) and peer discussion (Smith et al., 2009) in group settings should not be disregarded. Our findings showed that group settings can bridge understanding issues by boosting morale and letting participants share knowledge, interpretations, and experiences. We thus argue that individual and group settings support different understanding facets (Wiggins and McTighe, 2005), meaning that they provide different grounds for understanding AI systems. While individual settings can make it easier to understand technical and numerical details that require much attention (‘apply’), group settings can support understanding of the deployment context and consequences through the exchange of expertise and lived experiences (‘interpret’, ‘take perspective’, ‘empathize’). Consequently, individual and group settings should be combined to leverage their different modes of interaction and understanding facets when explaining AI systems. In cases where both social settings cannot be provided, explanations of AI systems should aim to reinforce facets that are not covered in the corresponding setting.
4.2. RQ2-Deliberation: How do AI novices use explanations to form opinions and make decisions about AI systems in groups and individual settings?
Before and after the explanation phase, participants decided if the study’s case should be deployed, and groups additionally made a collective decision (Table 4). We compared participants’ decisions and decision confidence and how they deliberated deployment to examine the impact of explanations and the social setting. We first describe participants’ confidence and decision changes in single interviews (4.2.1) and then present three cases of group discussion illustrating ‘elements of deliberation’ (Stromer-Galley, 2007), including reasoned arguments (4.2.2), disagreement (4.2.3), and groupthink (4.2.4). Section 4.2.5 summarizes the findings.
4.2.1. Explanation phase led to increased decision confidence and decision swings.
For most participants, deciding about the AMS algorithm’s deployment was a clear choice: 7 out of 8 groups and the majority of single participants voted “No” (Table 4). Many participants reported increased decision confidence after the explanation phase and stated that they felt better informed due to the explanations and, where applicable, the group discussion. Reasons for these increases included a better understanding of the system’s “fundamental idea” (S4) and the “exchange of different opinions and things that catch your eye” (G2). Participant B3 emphasized that the explanations, although they only contained factual information, provided a stark contrast to public narratives:
Well, I changed my mind – you think you understand something when you see it in the media. You have a political opinion about it. But you don’t know the background information. And when you get to the background information, you can have a completely different opinion. (B3)
This contrast highlights how explanations of an AI system can impact decision-making by correcting lay understandings (DeVito et al., 2018) and is in line with previous work that advocates for making algorithmic complexity visible to users to tackle lay understandings (Eslami et al., 2016). It further directly connects the explanations and changes in participants’ decision confidence. Few participants reported decreased decision confidence, and only S3 and S11 reported a stronger decrease of -2 (Table 4). S3 explained that the system might have benefits, but too much depended on the conditions for deployment. In particular, it should be deployed “responsibly, with a pilot project, in a selected group, for three months”, and not haphazardly, where “you sit around for a day or eight hours and then training is finished” (S3). S11, who together with S3 requested the most explanations out of all participants, paradoxically stated that their confidence decreased because “I’m still missing so much information. Especially […] how tedious it is for the counselors if they have to disagree with the system.” In consequence, S11 changed their decision from ‘Yes’ to ‘No’. Notably, this fear of an algorithmic imprint was a prevalent theme throughout all studies and was often connected to past experiences with digitization projects, and the corresponding institutional deficiencies. S5, who also changed from ‘Yes’ to ‘No’, similarly stated that the explanations helped them to scrutinize the system: “You don’t have to introduce anything that’s extra bad”. While the explanations thus made the decision more uncertain for some, they undoubtedly encouraged critical reflection about the use case and triggered decision changes. Despite only having “their own mind” (S8), single participants could make use of the explanations to weigh pros and cons and adjust their mental model. This form of “internal deliberation” is supported by exposure to different views, as provided through the argumentative explanations in usage and context. This suggests that explanations can substitute at least small parts of public deliberation, which is thought to be the more salient driver for “reasoning towards good outcomes” (Mercier and Landemore, 2012). To contrast these findings with focus groups, we illustrate this public deliberation with three conversation excerpts that showcase elements of deliberation in focus groups.
Focus groups | |||||||||
ID | Decision I | Group decision | Decision II | Decision Conf. | ID | Decision I | Group decision | Decision II | Decision Conf. |
A1 | No | Yes | 0 | F1 | Yes | Yes | 0 | ||
A2 | No | Yes | +1 | F2 | Yes | Yes | -1 | ||
A3 | Yes | Yes | +1 | F3 | Yes | No | +1 | ||
A4 | No | Yes | Yes | 0 | F4 | No | No | -1 | |
B1 | No | No | +3 | F5 | Yes | No | Yes | -1 | |
B2 | Yes | No | 0 | G1 | Yes | Yes | +1 | ||
B3 | Yes | No | +2 | G2 | No | No | 0 | ||
B4 | Yes | No | No | +3 | G3 | Yes | No | 0 | |
C1 | No | No | -1 | G4 | No | No | No | +1 | |
C2 | No | No | 0 | H1 | No | No | +1 | ||
C3 | No | No | Yes | +2 | H2 | No | No | +1 | |
D1 | No | No | +1 | H3 | No | No | +3 | ||
D2 | No | No | 0 | H4 | No | No | 0 | ||
D3 | No | No | Yes | 0 | H5 | No | No | No | +1 |
E1 | No | No | 0 | ||||||
E2 | No | No | 0 | ||||||
E3 | No | No | Yes | -1 | |||||
Single interviews | |||||||||
ID | Decision I | – | Decision II | Decision Conf. | ID | Decision I | – | Decision II | Decision Conf. |
S1 | Yes | – | Yes | 0 | S7 | No | – | No | +3 |
S2 | No | – | No | -1 | S8 | No | – | No | 0 |
S3 | No | – | No | -2 | S9 | No | – | No | +1 |
S4 | No | – | No | +1 | S10 | Yes | – | Yes | 0 |
S5 | Yes | – | No | +2 | S11 | Yes | – | No | -2 |
S6 | No | – | No | +2 | S12 | No | – | No | +1 |
4.2.2. Case 1 - reasoned arguments: Group B discusses whether to deploy the system.